find the y intercept and the x intercept for the equation 3x-2y=12
Answer by mananth(16949)

find the y intercept and the x intercept for the equation 3x-2y=12
...
put y = 0
3x-0=12
3x=12
/3
x=4 the x intercept
...
put x=0
0-2y=12
-2y=12
/-2
y=-6 the y intercept